# GUIDE: Enhancing Gradient Inversion Attacks in Federated Learning with Denoising Models

### v1

@@ 4.2 GUIDE Setup
In the setup phase, the attacker uses a surrogate dataset to train a denoising model, which will later be exploited in the reconstruction phase. This process is illustrated in phase (i) of Figure 2. Formally, let $\mathcal{A}(\nabla,f_{\theta})$ represent an attack procedure that, given the shared update $\nabla$ from the victim client, knowledge of the shared model $f_{\theta}$, and other basics information about local training procedure of clients (e.g., those described in Section 3) aims to reconstruct an approximation of the input data $\hat{X}=\{\hat{x}_{1},\ldots,\hat{x}_{N}\}$. The attack procedure $\mathcal{A}(\nabla,f_{\theta})$ operates iteratively, refining the estimates of the user’s inputs over $T$ iterations. According to the threat model described in Section 3, an attacker can exploit a surrogate dataset $D^{\prime}=\{(x^{\prime}_{i},y^{\prime}_{i})\}_{i=1}^{n^{\prime}}$ to create a set of noisy images through $\mathcal{A}$. In Algorithm 1, the attacker leverages the shared model and the surrogate dataset $D^{\prime}$ to compute a model update $\nabla^{\prime}$ over a batch of data $\{({x^{\prime}}_{i},y^{\prime}_{i})\}_{i=1}^{N}$ drawn from $D^{\prime}$ (line 5 of Algorithm 1). The model update, which depends on the learning algorithm used, may consist of gradients or model weights and is computed by the compute_model_update function.
@@
Then, the attacker applies the attack procedure $\mathcal{A}(\nabla^{\prime},f_{\theta})$ to obtain an estimated set of user inputs $\{\hat{x}_{i}\}_{i=1}^{N}$. As the proposed approach is designed to be independent of the specific attack $\mathcal{A}$, it is not possible to generalize when GUIDE can extract noisy images, as this is closely related to the behavior of $\mathcal{A}$. Therefore, we introduce $\mathcal{S}_{iters}$ in Algorithm 1 to formalize the set of iterations at which noisy images are sampled. In the case $\mathcal{S}_{iters}=\{T\}$, images are extracted only at the end of the attack procedure; otherwise, they are sampled during the attack process. By sampling images throughout intermediate iterations, rather than solely at the end of the attack, the attacker is enabled to capture and reconstruct images with different noise patterns at various stages of attack $\mathcal{A}$. This method enables the attacker to build a denoising dataset $D_{den}$, which consists of pairs of noisy images generated by the reconstruction attack and their corresponding original images from $D^{\prime}$. Using this dataset, the attacker trains a specialized denoising model $\mathcal{M}_{den}:\mathcal{X}\rightarrow\mathcal{X}$, intended to reduce noise in reconstructed images and improve fidelity. Details on the specific denoising model are provided in Section 4.4. By the end of the setup phase, the attacker tailors a denoising model ready-to-use for the reconstruction. In this work, the input space $\mathcal{X}\subseteq\mathbb{R}^{C\times W\times H}$ represents images with $C$ channels, width $W$, and height $H$. Although this study focuses on image data, according to most of the existing attacks, GUIDE is a general methodology potentially extensible to any GIAs in FL.
@@
4.3 GUIDE Reconstruction
This step involves the integration of the denoising model to enhance the reconstruction achieved by the attack procedure $\mathcal{A}$.
The attacker’s first step is to execute the setup phase, thereby obtaining a denoising model $\mathcal{M}_{den}$, which will be used to enhance reconstruction quality. Subsequently, the attack can be launched, choosing between the two configurations determined by the set $\mathcal{D}_{iters}$, which contains all iterations at which the denoising model must be applied.
In our work, the denoising model $\mathcal{M}_{den}$ is applied after the base reconstruction procedure $\mathcal{A}(\nabla,f_{\theta})$, i.e. when $\mathcal{D}_{iters}=\{T\}$. In this case, $\mathcal{A}(\nabla,f_{\theta})$ operates independently to generate an initial reconstruction of the victim client’s data, and $\mathcal{M}_{den}$ subsequently refines these noisy reconstructions $\hat{X}$ to produce cleaner versions.
@@ 4.3 GUIDE Reconstruction
ID Attack $\mathcal{A}$ Algorithm Task Model Training Client Data Surrogate Data Img. Resolution Denoising 1 CPA Kariyappa et al. (2023) FedSGD Img. Cls. VGG16 E=1E{=}1, $B{\in}\{4,\dots,256\}$ ImageNet (Test) Deng et al. (2009) ImageNet (Val) Deng et al. (2009) $224\times 224$ Cui et al. (2024) Luo et al. (2023b) Luo et al. (2023a) 2 ROG Yue et al. (2023) FedAVG Face Rec. ResNet-18 E=5E{=}5, B=16B{=}16 VGGFace2 (Subset) Cao et al. (2018) VGGFace2 (Non-Overlap IDs) Cao et al. (2018) $128\times 128$ Luo et al. (2023b) Luo et al. (2023a)
@@
Table 2: Overview of the attack scenarios and experimental setup, including image resolution.
4.4 Denoising Model
@@
Image denoising models are essential tools in image processing, aiming to recover high-quality images from their corrupted versions by reducing or eliminating noise. Traditional denoising methods often rely on the statistical properties of the noise and make assumptions about the underlying clean image structures. However, deep learning-based approaches have revolutionized the field by learning complex mappings from noisy to clean images directly from data, achieving state-of-the-art performance in various image restoration tasks Su et al. (2022). Among these models, those based on diffusion have recently outperformed other leading models in the field for this class of tasks Yang et al. (2023b); Li et al. (2023).
@@
Diffusion models employ a two-phase process: a forward and a reverse process. In the former, noise is gradually added to an image until it is indistinguishable from Gaussian noise. In the reverse process, the model gradually reduces noise by estimating and removing it at each step, progressively reconstructing the original image from a noisy input to achieve high-quality restoration. These models are usually evaluated in different configurations where the perturbations applied to the corrupted images follow specific patterns (e.g., Gaussian noise) Li et al. (2023).
@@
In our work, we leverage denoising models to enhance the quality of images reconstructed by existing attacks. Specifically, we consider the attack outputs $\mathcal{A}(\nabla,f_{\theta})$ as noisy versions of the client’s training images used to compute the gradient update $\nabla$. As mentioned in Section 3, the attacker can use a surrogate dataset $D^{\prime}$ to train the denoising model, which consists of images similar to those possessed by the clients in the FL setting. The noisy images used for training are generated using the attack $\mathcal{A}(\nabla,f_{\theta})$ itself, ensuring that the denoising model learns to invert the specific type of degradation introduced by the attack.
5 Experimental Analysis
@@
This section offers an extensive evaluation of GUIDE. First, we detail our experimental setup and investigate how the methodology can be evaluated across two attack scenarios using different FL algorithms, models, and datasets. Experimental results demonstrate that GUIDE remarkably enhances GIAs even in FL scenarios characterized by large batch sizes, high-dimensional data (i.e., 256 images of $\times224$ pixel) or when FedAVG is used as learning algorithm; conditions that typically reduce the efficacy of GIAs Du et al. (2025). Finally, we evaluate GUIDE’s resilience against defense mechanisms that may prevent the reconstruction, such as Differential Privacy (DP) Ponomareva et al. (2023) and heuristic gradient compression techniques like Quantized Stochastic Gradient Descent (QSGD) Alistarh et al. (2017) and Top-k Aji and Heafield (2017). All experiments were conducted using an NVIDIA A100-SXM4-80GB GPU within a high-performance infrastructure. For transparency and reproducibility, we will release our source code upon paper acceptance.
@@
Figure 3: Example of reconstructions obtained by different methods on 12 images in Scenario  1. The gradient is computed over 256 images with a resolution of $224\times 224$ with VGG16 model. All the images are not included in any noisy image seen during the training of the denoising models. (a) Images reconstructed with CPA Kariyappa et al. (2023). (b) Reconstruction with CPA Kariyappa et al. (2023) enhanced with GUIDE (ours). (c) Original images used to compute the gradient.
5.1 Experimental Setup
@@
In this section, we provide a detailed description of the experimental scenarios, evaluation methodology, and metrics used to assess how GUIDE improves reconstruction quality of GIAs.
@@
Attack Scenarios. To demonstrate the robustness and generalization of GUIDE in enhancing reconstruction quality, we consider two state-of-the-art attacks by evaluating its performance across two attack scenarios, whose main features are summarized in Table 2. These scenarios involve different attack strategies, shared models, and learning algorithms, allowing us to assess the methodology’s effectiveness under varying conditions. In both cases, the surrogate dataset used by the attacker is strictly non-overlapping with the private client’s dataset to ensure a fair and realistic assessment.
@@
In the first scenario, the adversary exploits CPA Kariyappa et al. (2023) for an image classification task using a VGG-16 Simonyan and Zisserman (2014) with FedSGD McMahan et al. (2017). In the second one, the attacker uses ROG Yue et al. (2023) for a face recognition task employing ResNet-18 He et al. (2016) with FedAVG McMahan et al. (2017). Both scenarios incorporate tailored denoising techniques to improve reconstruction quality. A detailed description of each scenario along with the denoising strategies employed and the result obtained is provided in Sections 5.2 and 5.3.
@@
Evaluation Metrics. In the evaluation of GIAs within the computer vision domain, image similarity is typically assessed using metrics such as Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) Zhang et al. (2018). PSNR measures the pixel-level fidelity between images, SSIM assesses structural similarities by considering luminance, contrast, and texture, while LPIPS leverages deep neural networks to approximate human perceptual judgments. However, these metrics primarily focus on pixel-wise differences and may not accurately reflect human perception when used to evaluate reconstructed images, particularly in terms of privacy leakage and recognizability Yue et al. (2023); Sun et al. (2024).
@@
To address these limitations, we also adopt the recently proposed DreamSim metric Fu et al. (2023), which is designed to better align with human perception. Unlike the above mentioned, DreamSim evaluates mid-level semantic variations, such as pose, perspective, and object shape, by leveraging high-level features from pre-trained vision models fine-tuned with human-aligned perceptual data. By computing perceptual similarity through cosine distance on feature embeddings, DreamSim offers a more comprehensive assessment of image recognizability, extending beyond pixel-level similarities. To the best of our knowledge, we are the first to use DreamSim for evaluating GIAs in FL, demonstrating its potential to offer a more accurate and human-aligned reconstruction quality assessment.
5.2 Attack Scenario 1: Image Classification
@@
Figure 4: (a) Examples of images generated by CPA Kariyappa et al. (2023) at various stages of the attack ($\mathcal{S}_{iters}=\{2000,8000,16000,20000,24000\}$) when gradients are computed over a batch of 256 images. (b) Reconstructed image obtained by starting from the image of (a) and applying denoising with a blended model. The denoising model demonstrates an ability to enhance reconstruction quality even in the early phases of the attack, as evidenced by lower DreamSim values.
@@
In this attack scenario, we consider CPA Kariyappa et al. (2023) as the base attack $\mathcal{A}$, utilizing a pre-trained VGG-16 Simonyan and Zisserman (2014) as the shared model. The model is trained with FedSGD for one local epoch, with batch sizes ranging from 4 and 256. The victim client’s dataset consists of $224\times 224$ images from the ImageNet Deng et al. (2009) test set, while the surrogate dataset is a subset of the validation set. To improve reconstruction quality, we employ denoising techniques based on both diffusion models Luo et al. (2023b, a) and convolutional denoising models Cui et al. (2024).
@@
GUIDE Setup Phase. In this phase, the attacker samples images from the surrogate dataset, computes model updates $\nabla^{\prime}$ using batch sizes ranging from 4 to 256, obtaining noisy images by applying CPA Kariyappa et al. (2023) as the base attack $\mathcal{A}$ to reconstruct the original data. In our experiments, the resulting noisy dataset $D_{den}$ consists of 106,008 pairs of images, each containing a noisy image and its corresponding original version. These images are collected at multiple iterations of 2000 of the CPA to train a denoising model capable of reconstructing images even during the intermediate iterations of the attack. A diffusion-based model and a convolutional model are employed to denoise the reconstructed images. After collecting the data, we performed an 80%/20% split for training and testing and trained both models using the original hyperparameters from Luo et al. (2023b); Cui et al. (2024). The diffusion model, based on a Conditional NAFNet architecture, was trained for 700,000 iterations with a batch size of 16, using the Lion optimizer Chen et al. (2024) and optimizing the $L_{1}$ loss function. The convolutional model, following the base architecture from Cui et al. (2024), was trained for 300 epochs with a batch size of 64 using the Adam Kingma (2014) optimizer. No data augmentation techniques were applied in either case. Based on the performance evaluation of both models, we selected a combined denoising approach as $\mathcal{M}_{den}$, which merges their outputs using a weighted fusion scheme (0.65 for the diffusion model and 0.35 for the convolutional model), as it yielded superior reconstruction quality.
@@
Table 3: Performance comparison of attacks when model update is computed across different batch sizes in Scenario  1. For each batch size, results are averaged over 30 different batches and reported as $mean_{\pm standard\;deviation}$. The best performance for each metric is highlighted in bold.
In Table 3, we report the results obtained by varying the batch size from 4 to 256 images. Across all batch sizes, GUIDE consistently achieves superior performance in terms of PSNR, with a mean improvement of 18%, and comparable results for SSIM, where GUIDE registers a mean improvement of 3%. Although the LPIPS metric favors CPA, a visual comparison (see Figure 3 and 4) reveals that images generated by CPA are not semantically closer to the originals. This suggests that LPIPS may not be an appropriate perceptual metric for evaluating privacy leakage in this context, as also discussed in Sun et al. (2024). Conversely, DreamSim highlights a substantial mean improvement of 27% for GUIDE over CPA. This result, further supported by visual inspection, indicates that DreamSim offers a more reliable assessment of privacy leakage in GIAs, in line with its proven effectiveness in other scenarios focused on perceptual similarity Fu et al. (2023); Ghazanfari et al. (2024).
@@
Table 4: Performance metrics for CPA Kariyappa et al. (2023) and GUIDE across different iterations. Each GUIDE configuration is evaluated over II iterations of the base CPA, with denoising applied only at the final iteration, denoted by $\mathcal{D}_{iters}=\{I\}$. Gradients are computed across 30 batches of 256 images, and results are averaged.
@@
Figure 4 clearly shows that GUIDE reconstructs images with significantly higher fidelity after 2,000 iterations compared to CPA. In this example, the visual findings is corroborated by a 69% improvement in the DreamSim metric when denoising is applied to the image obtained at iteration 2000. Additionally, Table 4 reports all the metrics when different numbers of iteration are used. With 2000 iterations, GUIDE achieves a higher PSNR than CPA, and this is higher as the number of iterations grows. By 8000 iterations, GUIDE reaches the highest PSNR value observed with an improvement of 15%. The DreamSim metric highlights the effectiveness of our approach as a comparable quality to CPA is achieved at the 2000-iteration mark, and superior reconstructions emerge from 4000 iterations onward. This suggests that GUIDE can yield more perceptually accurate images, even with a low number of iterations. This outcome indicates that attackers could achieve enhanced reconstruction quality with a reduced computational load. The computational burden can be shifted to the training phase of the denoising model, making it a flexible tool for attackers as it can be potentially deployed in multiple attack scenarios.
5.3 Attack Scenario 2: Face Recognition

### v2

@@ IV-B GUIDE Setup
In the setup phase, the attacker uses a surrogate dataset to train a denoising model, which will later be exploited in the reconstruction phase. This process is illustrated in phase (i) of Figure 2. Formally, let $\mathcal{A}(\nabla,f_{\theta})$ represent an attack procedure that, given the shared update $\nabla$ from the victim client, knowledge of the shared model $f_{\theta}$, and other basics information about local training procedure of clients (e.g., those described in Section III) aims to reconstruct an approximation of the input data $\hat{X}=\{\hat{x}_{1},\ldots,\hat{x}_{N}\}$. The attack procedure $\mathcal{A}(\nabla,f_{\theta})$ operates iteratively, refining the estimates of the user’s inputs over $T$ iterations. According to the threat model described in Section III, an attacker can exploit a surrogate dataset $D^{\prime}=\{(x^{\prime}_{i},y^{\prime}_{i})\}_{i=1}^{n^{\prime}}$ to create a set of noisy images through $\mathcal{A}$. In Algorithm 1, the attacker leverages the shared model and the surrogate dataset $D^{\prime}$ to compute a model update $\nabla^{\prime}$ over a batch of data $\{({x^{\prime}}_{i},y^{\prime}_{i})\}_{i=1}^{N}$ drawn from $D^{\prime}$ (line 5 of Algorithm 1). The model update, which depends on the learning algorithm used, may consist of gradients or model weights and is computed by the compute_model_update function.
@@
Then, the attacker applies the attack procedure $\mathcal{A}(\nabla^{\prime},f_{\theta})$ to obtain an estimated set of user inputs $\{\hat{x}_{i}\}_{i=1}^{N}$. As the proposed approach is designed to be independent of the specific attack $\mathcal{A}$, it is not possible to generalize when GUIDE can extract noisy images, as this is closely related to the behavior of $\mathcal{A}$. Therefore, we introduce $\mathcal{S}_{iters}$ in Algorithm 1 to formalize the set of iterations at which noisy images are sampled. In the case $\mathcal{S}_{iters}=\{T\}$, images are extracted only at the end of the attack procedure; otherwise, they are sampled during the attack process. By sampling images throughout intermediate iterations, rather than solely at the end of the attack, the attacker is enabled to capture and reconstruct images with different noise patterns at various stages of attack $\mathcal{A}$. This method enables the attacker to build a denoising dataset $D_{den}$, which consists of pairs of noisy images generated by the reconstruction attack and their corresponding original images from $D^{\prime}$. Using this dataset, the attacker trains a specialized denoising model $\mathcal{M}_{den}:\mathcal{X}\rightarrow\mathcal{X}$, intended to reduce noise in reconstructed images and improve fidelity. Details on the specific denoising model are provided in Section IV-D. By the end of the setup phase, the attacker tailors a denoising model ready-to-use for the reconstruction. In this work, the input space $\mathcal{X}\subseteq\mathbb{R}^{C\times W\times H}$ represents images with $C$ channels, width $W$, and height $H$. Although this study focuses on image data, according to most of the existing attacks, GUIDE is a general methodology potentially extensible to any GIAs in FL.
IV-C GUIDE Reconstruction
This stage involves the integration of the denoising model to enhance the reconstruction achieved by the attack procedure $\mathcal{A}$, as described in Algorithm 2. The attacker’s first step is to execute the setup phase, thereby obtaining a denoising model $\mathcal{M}_{den}$, which will be used to enhance reconstruction quality. Subsequently, the attack can be launched, choosing between the two configurations determined by the set $\mathcal{D}_{iters}$, which contains all iterations at which the denoising model must be applied. In our work, the denoising model $\mathcal{M}_{den}$ is applied after the base reconstruction procedure $\mathcal{A}(\nabla,f_{\theta})$, i.e. when $\mathcal{D}_{iters}=\{T\}$. In this case, $\mathcal{A}(\nabla,f_{\theta})$ operates independently to generate an initial reconstruction of the victim client’s data, and $\mathcal{M}_{den}$ subsequently refines these noisy reconstructions $\hat{X}$ to produce cleaner versions.
@@ IV-C GUIDE Reconstruction
ID Attack $\mathcal{A}$ Algorithm Task Model Training Client Data Surrogate Data Img. Resolution Denoising 1 CPA [20] FedSGD Img. Cls. VGG16 E=1E{=}1, $B{\in}\{4,\dots,256\}$ ImageNet [30] (Test) ImageNet [30] (Val) $224\times 224$ [31], [32], [33] 2 ROG [15] FedAVG Face Rec. ResNet-18 E=5E{=}5, B=16B{=}16 VGGFace2 [34] (Subset) VGGFace2 [34] (Non-Overlap IDs) $128\times 128$ [32], [33]
@@
TABLE II: Overview of the attack scenarios and experimental setup, including image resolution.
IV-D Denoising Model
@@
Image denoising models are essential tools in image processing, aiming to recover high-quality images from their corrupted versions by reducing or eliminating noise. Traditional denoising methods often rely on the statistical properties of the noise and make assumptions about the underlying clean image structures. However, deep learning-based approaches have revolutionized the field by learning complex mappings from noisy to clean images directly from data, achieving state-of-the-art performance in various image restoration tasks [35]. Among these models, those based on diffusion have recently outperformed other leading models in the field for this class of tasks [22, 36].
@@
Diffusion models employ a two-phase process: a forward and a reverse process. In the former, noise is gradually added to an image until it is indistinguishable from Gaussian noise. In the reverse process, the model gradually reduces noise by estimating and removing it at each step, progressively reconstructing the original image from a noisy input to achieve high-quality restoration. These models are usually evaluated in different configurations where the perturbations applied to the corrupted images follow specific patterns (e.g., Gaussian noise) [36].
@@
In our work, we leverage denoising models to enhance the quality of images reconstructed by existing attacks. Specifically, we consider the attack outputs $\mathcal{A}(\nabla,f_{\theta})$ as noisy versions of the client’s training images used to compute the gradient update $\nabla$. As mentioned in Section III, the attacker can use a surrogate dataset $D^{\prime}$ to train the denoising model, which consists of images similar to those possessed by the clients in the FL setting. The noisy images used for training are generated using the attack $\mathcal{A}(\nabla,f_{\theta})$ itself, ensuring that the denoising model learns to invert the specific type of degradation introduced by the attack.
V Experimental Analysis
@@
This section offers an extensive evaluation of GUIDE. First, we detail our experimental setup and investigate how the methodology can be evaluated across two attack scenarios using different FL algorithms, models, and datasets. Experimental results demonstrate that GUIDE remarkably enhances GIAs even in FL scenarios characterized by large batch sizes, high-dimensional data (i.e., 256 images of $\times224$ pixel) or when FedAVG is used as learning algorithm; conditions that typically reduce the efficacy of GIAs [9]. Finally, we evaluate GUIDE’s resilience against defense mechanisms that may prevent the reconstruction, such as Differential Privacy (DP) [37] and heuristic gradient compression techniques like Quantized Stochastic Gradient Descent (QSGD) [38] and Top-k [39]. All experiments were conducted using an NVIDIA A100-SXM4-80GB GPU within a high-performance infrastructure. For transparency and reproducibility, we will release our source code upon paper acceptance.
@@
Figure 3: Example of reconstructions obtained by different methods on 12 images in Scenario  1. The gradient is computed over 256 images with a resolution of $224\times 224$ with the VGG16 model. All the images are not included in any noisy image seen during the training of the denoising models. (a) Images reconstructed with CPA [20]. (b) Reconstruction with CPA [20] enhanced with GUIDE (ours). (c) Original images used to compute the gradient.
V-A Experimental Setup
@@
In this section, we provide a detailed description of the experimental scenarios, evaluation methodology, and metrics used to assess how GUIDE improves the reconstruction quality of GIAs.
@@
Attack Scenarios. To demonstrate the robustness and generalization of GUIDE in enhancing reconstruction quality, we consider two state-of-the-art attacks by evaluating its performance across two attack scenarios, whose main features are summarized in Table II. These scenarios involve different attack strategies, shared models, and learning algorithms, allowing us to assess the methodology’s effectiveness under varying conditions. In both cases, the surrogate dataset used by the attacker is strictly non-overlapping with the private client’s dataset to ensure a fair and realistic assessment.
@@
In the first scenario, the adversary exploits CPA [20] for an image classification task using a VGG-16 [40] with FedSGD [2]. In the second one, the attacker uses ROG [15] for a face recognition task employing ResNet-18 [41] with FedAVG [2]. Both scenarios incorporate tailored denoising techniques to improve reconstruction quality. A detailed description of each scenario, along with the denoising strategies employed and the result obtained, is provided in Sections V-B and V-C.
@@
Evaluation Metrics. In the evaluation of GIAs within the computer vision domain, image similarity is typically assessed using metrics such as Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) [42]. PSNR measures the pixel-level fidelity between images, SSIM assesses structural similarities by considering luminance, contrast, and texture, while LPIPS leverages deep neural networks to approximate human perceptual judgments. However, these metrics primarily focus on pixel-wise differences and may not accurately reflect human perception when used to evaluate reconstructed images, particularly in terms of privacy leakage and recognizability [15, 23].
@@
To address these limitations, we also adopt the recently proposed DreamSim metric [24], which is designed to better align with human perception. Unlike the above-mentioned, DreamSim evaluates mid-level semantic variations, such as pose, perspective, and object shape, by leveraging high-level features from pre-trained vision models fine-tuned with human-aligned perceptual data. By computing perceptual similarity through cosine distance on feature embeddings, DreamSim offers a more comprehensive assessment of image recognizability, extending beyond pixel-level similarities. To the best of our knowledge, we are the first to use DreamSim for evaluating GIAs in FL, demonstrating its potential to offer a more accurate and human-aligned reconstruction quality assessment.
V-B Attack Scenario 1: Image Classification
@@
Figure 4: (a) Examples of images generated by CPA [20] at various stages of the attack ($\mathcal{S}_{iters}=\{2000,8000,16000,20000,24000\}$) when gradients are computed over a batch of 256 images. (b) Reconstructed image obtained by starting from the image of (a) and applying denoising with a blended model. The denoising model demonstrates an ability to enhance reconstruction quality even in the early phases of the attack, as evidenced by lower DreamSim values.
@@
In this attack scenario, we consider CPA [20] as the base attack $\mathcal{A}$, utilizing a pre-trained VGG-16 [40] as the shared model. The model is trained with FedSGD for one local epoch, with batch sizes ranging from 4 and 256. The victim client’s dataset consists of $224\times 224$ images from the ImageNet [30] test set, while the surrogate dataset is a subset of the validation set. To improve reconstruction quality, we employ denoising techniques based on both diffusion models [32, 33] and convolutional denoising models [31].
@@
GUIDE Setup Phase. In this phase, the attacker samples images from the surrogate dataset, computes model updates $\nabla^{\prime}$ using batch sizes ranging from 4 to 256, obtaining noisy images by applying CPA [20] as the base attack $\mathcal{A}$ to reconstruct the original data. In our experiments, the resulting noisy dataset $D_{den}$ consists of 106,008 pairs of images, each containing a noisy image and its corresponding original version. These images are collected at multiple iterations of 2000 of the CPA to train a denoising model capable of reconstructing images even during the intermediate iterations of the attack. A diffusion-based model and a convolutional model are employed to denoise the reconstructed images. After collecting the data, we performed an 80%/20% split for training and testing and trained both models using the original hyperparameters from [32, 31]. The diffusion model, based on a Conditional NAFNet architecture, was trained for 700,000 iterations with a batch size of 16, using the Lion optimizer [43] and optimizing the $L_{1}$ loss function. The convolutional model, following the base architecture from [31], was trained for 300 epochs with a batch size of 64 using the Adam [44] optimizer. No data augmentation techniques were applied in either case. Based on the performance evaluation of both models, we selected a combined denoising approach as $\mathcal{M}_{den}$, which merges their outputs using a weighted fusion scheme (0.65 for the diffusion model and 0.35 for the convolutional model), as it yielded superior reconstruction quality.
@@
TABLE III: Performance comparison of attacks when model update is computed across different batch sizes in Scenario  1. For each batch size, results are averaged over 30 different batches and reported as $mean_{\pm standard\;deviation}$. The best performance for each metric is highlighted in bold.
In Table III, we report the results obtained by varying the batch size from 4 to 256 images. Across all batch sizes, GUIDE consistently achieves superior performance in terms of PSNR, with a mean improvement of 18%, and comparable results for SSIM, where GUIDE registers a mean improvement of 3%. Although the LPIPS metric favors CPA, a visual comparison (see Figure 3 and 4) reveals that images generated by CPA are not semantically closer to the originals. This suggests that LPIPS may not be an appropriate perceptual metric for evaluating privacy leakage in this context, as also discussed in [23]. Conversely, DreamSim highlights a substantial mean improvement of 27% for GUIDE over CPA. This result, further supported by visual inspection, indicates that DreamSim offers a more reliable assessment of privacy leakage in GIAs, in line with its proven effectiveness in other scenarios focused on perceptual similarity [24, 25].
@@
TABLE IV: Performance metrics for CPA [20] and GUIDE across different iterations. Each GUIDE configuration is evaluated over II iterations of the base CPA, with denoising applied only at the final iteration, denoted by $\mathcal{D}_{iters}=\{I\}$. Gradients are computed across 30 batches of 256 images, and results are averaged.
@@
Figure 4 clearly shows that GUIDE reconstructs images with significantly higher fidelity after 2,000 iterations compared to CPA. In this example, the visual findings are corroborated by a 69% improvement in the DreamSim metric when denoising is applied to the image obtained at iteration 2000. Additionally, Table IV reports all the metrics when different numbers of iterations are used. With 2000 iterations, GUIDE achieves a higher PSNR than CPA, and this is higher as the number of iterations grows. By 8000 iterations, GUIDE reaches the highest PSNR value observed with an improvement of 15%. The DreamSim metric highlights the effectiveness of our approach as a comparable quality to CPA is achieved at the 2000-iteration mark, and superior reconstructions emerge from 4000 iterations onward. This suggests that GUIDE can yield more perceptually accurate images, even with a low number of iterations. This outcome indicates that attackers could achieve enhanced reconstruction quality with a reduced computational load. The computational burden can be shifted to the training phase of the denoising model, making it a flexible tool for attackers as it can be potentially deployed in multiple attack scenarios.
V-C Attack Scenario 2: Face Recognition